The screen size of a t.V is the length of the diagonal across the screen. The height of the screen = 16inches. The diagonal of the screen is 32inches. How to i find out the width?
Answer by actuary(112)

This problem is solved by using the Pythagrean Theorem. The diagonal of the TV screen is the hypotenuse of a right triangle with the sides (legs) of the triangle formed by the sides of the TV screen.
So 32^2 = 16^2 + w^2 (Pythagorean Theorem a^2 + b^2 = c^2)
So 1024 = 256 + w^2 and then w^2 = 1024 - 256 = 768
Therefore, w = Square Root [768] = 27.71 in.